40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 9 15:49:24 2011, Jenne, Update, IOO, MC trans is way down MCtrans_9Dec2011.png
    Reply  Fri Dec 9 17:15:29 2011, Jenne, Update, IOO, MC trans is way down 
Message ID: 6098     Entry time: Fri Dec 9 17:15:29 2011     In reply to: 6096
Author: Jenne 
Type: Update 
Category: IOO 
Subject: MC trans is way down 

Quote:

I was looking at the trends from the RAMmon, since I did the EOM alignment yesterday, and wanted to compare them to the MC trans, just to make sure the MC was locked during the times I'm examining.  I was dismayed to discover that the MC has lost its oomph, starting around 11:30 this morning.  Den was the only person in the lab to my knowledge at that time, and he claims that he didn't touch the MC until well after lunch.  As you can see from the 8 hour trend attached, we went from normal ~26000 counts to ~15000, and we're slowly decaying from there.

MC refl looks pretty bad on the camera, particularly in YAW.  Investigations are beginning now....

Edit, ~10min later...  I enabled the WFS (I don't know why they were off...when the MC fell out of lock and relocked itself, the WFS didn't come on), and things went basically back to how they should be.  However, the sans-WFS alignment is still totally crappy, so the PSL beam probably needs to be aligned to the MC.  I don't really want to touch the alignment though without an okay from Kiwamu, so I'll wait for him to come in and confirm that he's happy with the current MC.

 Kiwamu and I discussed, and looking at the AS camera with the PRM and SRM misaligned, but MCWFS engaged, things look good.  This means that it's probably the MC that has drifted, and we want to align the MC back to the PSL beam.

ELOG V3.1.3-